# The EMPI Code for Plasma-Induced Effects on Radio Waves I: Non-Magnetized Media and Applications to Fast Radio Bursts

### v1

@@ 4 Results
This section presents numerical results, organized into two parts. The first part focuses on code validation by comparing numerical outcomes with analytical solutions. The second part explores scenarios beyond the scope of analytical descriptions, demonstrating the capabilities of the code in handling complex situations. Throughout the section, the input light curve and spectrum remain fixed. A rectangular pulse with a duration of $2$ ms is used for the lightcurve and the spectrum is a flat-top rectangular wave spanning the frequency range of 1–1.5 GHz.
@@ 4.1 Validation of Standard Results
This section rigorously evaluates the resolution and capabilities of this code. Its performance is systematically validated by comparing computational results against analytical solutions across multiple benchmarks. First, the ability of this code to reproduce four non-magnetized plasma-induced effects—dispersion, plasma lensing, scintillation, and scattering—is demonstrated, establishing an initial overview of its standard outputs. Subsequently, the numerical resolution of the code is assessed to reinforce confidence in its reliability.
@@ 4.1.1 Dispersion Test
The dispersion effect, a ubiquitous phenomenon in the presence of plasma, manifests as the time delay between the low- and high-frequency components of the signal. This delay, termed dispersion delay, is defined as:
@@
where $\nu_{2}>\nu_{1}$ and the condition $\omega\gg\omega_{p}$ holds (Zhang, 2023). To isolate and validate the dispersion effect in the code, other plasma-induced phenomena (lensing, scintillation, scattering) are excluded. The light is set to propagate through a uniform plasma screen with a density of $20\ {\rm cm^{-3}}$ and a thickness of $0.1\ {\rm pc}$. Since no deflection occurs, only light traveling along the $x$-axis (from the source to the observer) is received. Simulated results are shown in panel (a) of Figure 5, displaying the raw signal waterfall plot without dedispersion. This plot acts as a reference for the standard output of this code, annotated with the true DM value and the resolutions of the telescope. The true DM (calculated numerically as the average DM) and its standard deviation are derived directly from the traced light rays, each contributing a unique DM value during superposition at the observer.
@@
The simulated signals closely mimic real observational received signals, enabling analogous processing to derive physical insights. For instance, the dispersed signal shown in the left panel of Figure 5, can be dedispersed to estimate a fitted DM value. The dedispersed result, displayed in the right panel of Figure 5, demonstrates the recovery of a near-flat-top lightcurve. The optimal DM is determined by iteratively shifting the signal across trial DM values and identifying the value that maximizes the peak flux. [^3] The dedispersion method iterates over a grid of trial DM values spanning the plausible range of DM values. For each trial DM value, frequency-dependent time delays are calculated using Eq. 13 and applied inversely to align frequency channels. The aligned data are integrated to generate trial lightcurves, with the optimal DM identified as the value maximizing the peak flux. This approach is consistent with the incoherent dedispersion method described in the review by Petroff et al. (2019). This yields a fitted DM of $1.96\ \ {\rm pc\ cm^{-3}}$, consistent with the true physical value of $2.00\ \ {\rm pc\ cm^{-3}}$ (relative error: $2\%$). While the dedispersed lightcurve approximates a rectangular pulse, its edges exhibit slight curvature rather than sharp transitions. This imperfection arises from finite temporal resolution: edge time bins contain partial signal contributions, compounded by residual inaccuracies in the dedispersion.
@@ 4.1.2 Plasma Lensing Test
Numerically, plasma lensing is modeled in the refractive regime by iteratively updating light-ray deflection angles at each propagation step. To validate this implementation, the following test evaluates the accuracy of predicted light deflection.
After constructing the plasma density grid, the Light Propagation Simulator computes light-ray trajectories through the medium (see Sections 2.2 and 3.2), resolving deflection angles, time delays, and phase shifts. Although full trajectory reconstruction remains intractable for explicit visualization, intersection points at the observer plane are captured through parametric coordinate mapping. Figure 6 illustrates typical arrival point distributions under refractive conditions. Panel (a) corresponds to a uniform plasma screen, where no deflection occurs. Intersection points remain frequency-independent and uniformly distributed across the observer plane. Because of the complete spatial overlap of coordinates across frequencies, the lower-frequency components become invisible in the visualization, leaving only the highest frequency remaining in the plot. However, observational measurements still preserve distinct spectral signatures. In contrast, Panel (b) depicts a divergent plasma lens. Lower frequencies experience stronger deflections, spreading light over a broader area. Incident parallel light that is initially confined to a $10$ AU radius circular region on the plasma screen. Since the incident light is parallel the observer plane also has a circle of 10 AU radius on the observer plane for the uniform case (Panel (a)). For the divergent lens (Panel (b)), the light disperse across $\sim-1000$ AU to $\sim 1000$ AU, starkly contrasting with the uniform screen and confirming the lensing divergence.
@@
By dividing the observer plane into discrete bins and quantifying the density of intersection points within each bin, the relative gain of the plasma lens is derived. A higher density of points corresponds to elevated intensity levels, as illustrated in Figure 7. Panels (a) and (b) of Figure 7 correspond to their counterparts in Figure 6, representing uniform and divergent plasma configurations, respectively. Panel (a) represents the uniform plasma case, no deflection occurs, resulting in the uniform gain across the observer plane. The edge decline arises from incomplete sampling in peripheral bins, where fewer rays contribute. In contrast, Panel (b) shows the divergent lens case that redistributes light non-uniformly, broadening the spatial spread due to frequency-dependent deflections. The divergent lens produces a frequency-dependent intensity distribution, concentrating light into caustic structures while depleting flux in the central region. Caustic positions shift systematically with frequency, reflecting the chromatic nature of plasma lensing.  [^4] Panel (b) may appear to exhibit light concentration; however, this is an artifact of the differing $y$- and $z$-axis scales between panels. The apparent bright regions arise from caustics formed by deflected light, rather than the convergent lens effect.
@@
To analyze the received signal at a specific observer-plane location, one may tally light rays across all frequencies arriving at that point. However, finite numerical resolution may result in few or no rays reaching the selected coordinate (see Section 4.1.5 for detailed discussion). Global grid refinement across the entire plane is computationally wasteful for single-point analyses. Instead, the region of interest and its immediate vicinity can be targeted, as demonstrated in Figure 8 (a). By tracing the path of each light ray, the corresponding screen patch passed through by the ray is identified. The code automatically isolates the relevant screen regions (as shown in Figure 8 (b)), allowing the targeted area to be divided solely into finer grids with higher resolution.
@@
Since the calculation of light ray deflection is a central component of the code, its numerical results require rigorous validation through quantitative analysis. The deflection accuracy is verified by benchmarking against analytical solutions from plasma lensing literature. Prior works, including studies by (Clegg et al., 1998) and (Cordes et al., 2017), model Gaussian plasma lenses as one-dimensional (1D) systems. To align with these models, the plasma screen is configured with no variation along the $x$-axis and $z$-axis, effectively simplifying the scenario to a 1D case. This enables direct comparison of the 1D numerical results with published analytical predictions.
In a one-dimensional plasma lens system, the fundamental quantity is the electron column density distribution, which, for a Gaussian profile, is expressed as:
@@
The lensing effects depend not solely on this distribution but also on additional physical parameters, and certain combinations of these parameters exhibit degeneracy, meaning that distinct sets of input parameters can result in observationally indistinguishable lensing outcomes. To facilitate quantitative characterization of the lensing effects, a dimensionless parameter $\alpha$ is introduced, which is recognized in current research as a fundamental descriptor of the lens characteristics. This critical parameter is defined (Clegg et al., 1998):
@@
To validate the methodology, numerical parameters are calibrated to correspond with the configuration in Clegg et al. (1998), with comparative results presented in Figure 9. The dimensionless intensity $I/I_{\rm 0}$ (where $I_{\rm 0}$ denotes unlensed uniform intensity) is plotted versus the normalized transverse coordinate $u=y_{\rm obs}/a$, representing the observer plane. Numerical results (red dotted curves) demonstrate excellent agreement with analytical solutions (blue solid curves), exhibiting only minor numerical fluctuations. Quantitative error analysis was performed through grid-matched interpolation, calculating pointwise relative error $|(I_{\rm num}-I_{\rm anal})/I_{\rm anal}|$. Mean relative errors across the profiles measure $0.3\%$, $0.5\%$, $9.2\%$, $10.3\%$, $3.9\%$ and $7.3\%$ for panels (a)-(f) respectively. Notably, elevated errors in panels (c)-(f) originate from caustic regions, where analytical solutions diverge at singularities. Since numerical values cannot technically be divergent, significant errors occur near the critical divergent points. However, as shown in the figure, the numerical results maintains consistence with the analytical ones.
@@ 4.1.3 Scintillation Test
Scintillation constitutes a fundamental plasma propagation effect in non-magnetized media, characterized by the variations of intensities across temporal and spectral domains. This phenomenon originates from coherent electromagnetic wave interference through multi-path propagation induced by plasma density fluctuations. As radiation traverses inhomogeneous plasma structures, differential phase accumulation along distinct paths generates interference patterns when light components reach the observer. In observations, scintillation manifests as characteristic intensity modulation patterns in dynamic spectra, exhibiting alternating bright and dark striations in frequency-time space.
In this framework, scintillation is fundamentally associated with the Light Receiver module, and its accurate reproduction validates the numerical implementation of coherent signal summation—a critical component of the code. Given the intrinsic coherence of FRB signals, interference effects require rigorous treatment via the phased summation methodology detailed in Section 2.5. The phase difference, which is essential in determining whether interference leads to amplitude enhancement or cancellation, arises from both geometric path variations and dispersive effects due to density fluctuations (i.e., dispersion measure variations along the paths). Although conventional scintillation theories predominantly address geometric phase contributions, this code self-consistently incorporates both geometric and dispersive phase components. To verify that the code properly handles phase differences and coherent signal superposition, an example considering only geometric phase shifts is provided. In this configuration, an infinitely distant source and two patches (or slits) allow light to pass through: One patch lies along the source-observer axis, and the other is parallel to the first but laterally offset by a distance $x$ within the plasma plane. Both patches are located at a distance $D$ from the observer plane. This arrangement mirrors the classic two-slit interference experiment; however, while the traditional two-slit experiment focuses on the formation of bright and dark fringes at a fixed frequency on the observer plane, scintillation refers to the phenomenon in which an observer at a fixed location receives light spanning different frequencies. This configuration generates frequency-dependent interference through the geometric phase delay:
@@
Consequently, the phase differences between the light from the two patches vary with frequency, leading to constructive interference at some frequencies and destructive interference at others, thereby producing scintillation patterns that appear as stripes in the waterfall plot of the received signal. One direct approach to estimate the scintillation bandwidth is outlined here, wherein the difference of phase differences between different frequencies can be expressed as:
@@
[…]
@@ 4.1.4 Scattering Test
Scattering, the last effect of non-magnetized plasma examined in this study, manifests primarily as time broadening, an effect that is particularly pronounced at low frequencies. In the literature, this phenomenon is attributed to a scattering screen, which is modeled in the diffractive regime in the EMPI code. Only light rays within a specific effective radius defined by the scattering screen can reach the observer. Because these rays follow varying geometric trajectories from different points within the effective circle, their arrival times differ, resulting in an extended signal duration that appears as a scattering tail in the waterfall plot. At lower frequencies, the effective radius increases, leading to a wider range of arrival times and, consequently, more pronounced signal broadening. This characteristic tail-like feature remains evident even after dedispersion, as illustrated in Figure 11 (the same figure also reveals a rising wing preceding the main pulse flattening; this is due to the intrinsic pulse being a top-flat square wave, deliberately chosen to accentuate these effects, with both the scattering tail and the rising wing being influenced by the dedispersion process).
@@
4.1.5 Resolution Test
For completeness, a resolution test is presented in this section. Numerical calculations inherently require subdividing the problem domain into small units, a process that may introduce discontinuities and fluctuations in the results. In general, as the resolution increases, the results converge toward a stable value that is considered to represent the final outcome.
The code under investigation produces two types of outputs — the simulated received signal and the plasma lensing gain — both of which are influenced by the resolution of the screen. The following demonstrates how these outputs converge to stable values as the resolution is progressively refined.

### v2

@@ 4 Results
This section presents numerical results, organized into two parts. The first part focuses on code validation by comparing numerical outcomes with analytical solutions. The second part explores scenarios beyond the scope of analytical descriptions, demonstrating the capabilities of the code in handling complex situations. Throughout the section, the input lightcurve and spectrum remain fixed. A rectangular pulse with a duration of $2$ ms is used for the lightcurve and a flat-top rectangular spectrum spans $1-1.5$ GHz.
@@ 4.1 Validation of Standard Results
This section rigorously evaluates the capabilities and convergence of this code. Its performance is validated through systematic comparison of computational results against analytical solutions across multiple benchmarks. First, the ability of this code to reproduce four nonmagnetized plasma-induced effects—dispersion, plasma lensing, scintillation, and scattering—is demonstrated, establishing a baseline for its standard outputs. Subsequently, numerical convergence is assessed by incrementally increasing the grid resolution to validate its reliability.
@@ 4.1.1 Dispersion Test
The dispersion effect, a ubiquitous phenomenon in the presence of plasma, manifests as the time delay between the low- and high-frequency components of the signal. This delay, termed dispersion delay, is defined as
@@
where $\nu_{2}>\nu_{1}$ and the condition $\omega\gg\omega_{p}$ holds (Zhang, 2023). To isolate and validate the dispersion effect in the code, other plasma-induced phenomena (lensing, scintillation, scattering) are excluded. The light is set to propagate through a uniform plasma screen with a density of $20\ {\rm cm^{-3}}$ and a thickness of $0.1\ {\rm pc}$. Since no deflection occurs, only light traveling along the $x$-axis (from the source to the observer) is received. Simulated results are shown in panel (a) of Figure 5, displaying the raw signal waterfall plot without dedispersion. This plot acts as a reference for the standard output of this code, annotated with the true DM value and the resolutions of the telescope. The true DM (statistical average of individual DM values from all rays reaching the observer) and its standard deviation are derived directly from the ray-tracing results. Each received ray contributes a unique DM value during superposition at the observer.
@@
The simulated signals achieve high fidelity to real observational data, thereby permitting application of analogous processing methodologies to extract physical insights. For instance, the dispersed signal shown in the left panel of Figure 5, can be dedispersed to estimate a fitted DM value. The dedispersed result, displayed in the right panel of Figure 5, demonstrates the recovery of a near-flat-top lightcurve. The optimal DM is determined by iteratively shifting the signal across trial DM values and identifying the value that maximizes the peak flux. [^3] The dedispersion method iterates over a grid of trial DM values spanning the plausible range of DM values. For each trial DM value, frequency-dependent time delays are calculated using Eq. 13 and applied inversely to align frequency channels. The aligned data are integrated to generate trial lightcurves, with the optimal DM identified as the value maximizing the peak flux. This approach is consistent with the incoherent dedispersion method described in the review by Petroff et al. (2019). This yields a fitted DM of $1.96\ \ {\rm pc\ cm^{-3}}$, consistent with the true physical value of $2.00\ \ {\rm pc\ cm^{-3}}$ (relative error: $2\%$). While the dedispersed lightcurve approximates a rectangular pulse, its edges exhibit slight curvature rather than sharp transitions. This imperfection stems from finite temporal resolution: edge time bins capture incomplete signal segments, further exacerbated by residual inaccuracies in the dedispersion process.
@@ 4.1.2 Plasma Lensing Test
Computationally, plasma lensing is modeled within the refractive regime through iterative recalculation of light-ray deflection angles at each propagation step. To validate this implementation, the following test evaluates the accuracy of predicted light deflection.
After constructing the plasma density grid, the light propagation simulator computes light-ray trajectories through the medium (see Sections 2.2 and 3.2), resolving deflection angles, time delays, and phase shifts. Although full trajectory reconstruction is impractical for explicit visualization, intersection points at the observer plane are captured via parametric coordinate mapping. Figure 6 illustrates typical arrival point distributions. Panel (a) corresponds to a uniform plasma screen, where no deflection occurs. Intersection points are frequency independent and uniformly distribute across the observer plane. Because of the complete spatial overlap across frequencies, lower-frequency components become invisible in the visualization (only the highest frequency remaining visible). However, this is purely a graphical limitation–no physical obscuration occurs and observational data retains full spectral information. Incident parallel light that is initially confined to a $10$ au radius circular region on the plasma screen. Since the incident light is parallel, the observer plane also has a circle of $10$ au radius on the observer plane for the uniform case (panel (a)). In contrast, panel (b) depicts a divergent plasma lens. Lower frequencies experience stronger deflections, spreading light over a broader area. The light disperses across $\sim 1000$ au to $\sim 1000$ au, starkly contrasting with the uniform screen and confirming the lensing divergence.
@@
By dividing the observer plane into discrete bins and statistically analyzing the density of intersection points within each bin, the relative gain induced by the plasma lens can be quantified. As illustrated in Figure 7, regions with higher intersection densities correspond to amplified intensity levels. Panels (a) and (b) of Figure 7 correspond to their counterparts in Figure 6, representing uniform and divergent plasma configurations, respectively. Panel (a) represents the uniform plasma case where no deflection occurs, resulting in uniform gain across the observer plane. The apparent intensity decline at the edges is an artifact of incomplete sampling in peripheral bins, where fewer rays contribute. By contrast, panel (b) shows the divergent lens case that redistributes light nonuniformly, broadening the spatial spread through chromatic deflections. The divergent lens produces a frequency-dependent intensity distribution, concentrating light into caustic structures while depleting flux in the central region. Caustic positions shift systematically with frequency, reflecting the chromatic nature of plasma lensing [^4] Panel (b) may appear to exhibit light concentration; however, this is an artifact of the differing $y$- and $z$-axis scales between panels. The apparent bright regions arise from caustics formed by deflected light, rather than the convergent lens effect..
@@
To analyze the received signal at a specific observer plane location, one may tally light rays across all frequencies arriving at that point. However, finite numerical resolution may result in few or no rays reaching the selected coordinate (see Section 4.1.5 for detailed discussion). Global grid refinement across the entire screen is computationally wasteful for single point analyses. Instead, the region of interest and its immediate vicinity can be targeted, as demonstrated in Figure 8 (a). By tracing individual light trajectories, the code identifies the specific screen patches passed through by each ray. This automated localization of relevant screen regions (as shown in Figure 8 (b)), enables selective subdivision of the targeted area into high-resolution fine grids.
@@
As the computation of light-ray deflection constitutes the core algorithm of the code, its numerical results necessitates rigorous validation through quantitative benchmarking. The deflection calculations are verified against analytical solutions for plasma lensing in prior theoretical studies. Previous works, including studies by (Clegg et al., 1998) and (Cordes et al., 2017), model Gaussian plasma lenses as one-dimensional (1D) systems. To align with these models, the plasma screen is configured with no variation along the $x$-axis and $z$-axis, effectively simplifying the scenario to a 1D case. This enables direct comparison of the 1D numerical results with published analytical predictions.
In a 1D plasma lens system, the fundamental quantity is the electron column density distribution, which, for a Gaussian profile, is expressed as
@@
The lensing effects depend not only on the plasma distribution but also on additional physical parameters. Crucially, parameter degeneracy arises when distinct combinations of these inputs yield observationally indistinguishable lensing signatures. To facilitate quantitative characterization of the lensing effects, a dimensionless parameter $\alpha$ is introduced, which is recognized in current research as a fundamental descriptor of the lens characteristics. This critical parameter is defined (Clegg et al., 1998) as
@@
To validate the methodology, numerical parameters are calibrated to correspond with the configuration in Clegg et al. (1998), with comparative results presented in Figure 9. The dimensionless intensity $I/I_{\rm 0}$ (where $I_{\rm 0}$ denotes unlensed uniform intensity) is plotted against the normalized transverse coordinate $u=y_{\rm obs}/a$, parameterizing positions on the observer plane. Numerical results (red dotted curves) demonstrate excellent agreement with analytical solutions (blue solid curves), exhibiting only minor numerical fluctuations. Quantitative error analysis employs grid-matched interpolation to compute pointwise relative error $|(I_{\rm num}-I_{\rm anal})/I_{\rm anal}|$. Mean relative errors across the profiles measure $0.3\%$, $0.5\%$, $9.2\%$, $10.3\%$, $3.9\%$ and $7.3\%$ for panels (a)-(f) respectively. Elevated errors in panels (c)–(f) originate from caustic regions, where analytical solutions diverge at singularities. Since numerical values cannot technically be divergent, significant errors occur near the critical divergent points. Crucially as shown in the figure, the numerical results maintain consistent with the analytical ones.
@@ 4.1.3 Scintillation Test
Scintillation constitutes a fundamental plasma propagation effect in nonmagnetized media, characterized by the variations of intensities across temporal and spectral domains. This phenomenon originates from coherent electromagnetic wave interference through multipath propagation induced by plasma density fluctuations. As radiation traverses inhomogeneous plasma structures, different phase accumulation along distinct paths generates interference patterns when light components reach the observer. In observations, scintillation manifests as characteristic intensity modulation patterns in dynamic spectra, exhibiting alternating bright and dark stripes in the frequency-time space.
In this framework, scintillation is inherently associated with the light receiver module, and its accurate reproduction validates the numerical implementation of coherent signal summation–the core functionality enabling interference modeling. Given the intrinsic coherence of FRB signals, interference effects demand rigorous treatment via the phased summation methodology detailed in Section  2.5. The phase difference, which determines whether interference leads to amplitude enhancement or cancellation, arises from both geometric path variations and dispersive effects due to electron density fluctuations (i.e., DM variations along the paths). Although conventional scintillation theories predominantly address geometric phase contributions, this code self-consistently incorporates both geometric and dispersive phase components. To verify that the code properly handles phase differences and coherent signal superposition, an example considering only geometric phase shifts is provided. In this configuration, two patches (or slits) allow light from an infinitely distant source to pass through. One patch lies along the source-observer axis, and the other is parallel to the first but laterally offset by a distance $x$ within the plasma plane. Both patches are located at a distance $D$ from the observer plane. This arrangement mirrors the classic two-slit interference experiment, yet diverges fundamentally in its observational manifestation: while the traditional two-slit experiment focuses on the formation of bright and dark fringes at a fixed frequency on the observer plane, scintillation refers to the phenomenon in which an observer at a fixed location receives light spanning different frequencies. This configuration generates frequency-dependent interference through the geometric phase delay:
@@
Consequently, the phase differences between the light from the two patches vary with frequency, leading to constructive interference at some frequencies and destructive interference at others, thereby producing scintillation patterns that appear as stripes in the waterfall plot of the received signal. One direct approach to estimate the scintillation bandwidth is formulated below, where the difference of phase differences between different frequencies can be expressed as
@@
Varied values of $D$ are applied in the panels of Figure 10 to demonstrate the ability of the code to effectively handle phase differences and coherent signal superposition. In this setup, we fix $x=10^{13}$ cm and choose distances $D$ of $1$ kpc, $10$ kpc, and $100$ kpc for Panels (a), (b), and (c), respectively. The estimated scintillation bandwidths from Eq. 18 are $0.93$ MHz, $9.3$ MHz, and $93$ MHz. Fourier transformation yields corresponding bandwidths of $1.1$ MHz, $9.1$ MHz, and $100$ MHz, which align well with the analytical predictions, with relative errors of $18\%$, $2\%$, and $7.5\%$, respectively. The elevated errors in cases (a) and (c) can be attributed to specific factors. When implementing a $1$ MHz spectral resolution, case (a) exhibits pronounced beat patterns caused by interference between the instrument resolution and the scintillation bandwidth. To mitigate this effect, panel (a) employs a spectral resolution of $0.5$ MHz, whereas Panels (b) and (c) maintain a resolution of $1$ MHz. Nonetheless, residual beat modulation persists, remaining higher relative errors in panel (a). In panel (c), the approximate $100$ MHz scintillation bandwidth occupies half of the $200$ MHz observational bandwidth, limiting the effective sampling window for Fourier transformation, which further contributes to the elevated relative error.
@@ 4.1.4 Scattering Test
Scattering, the final effect of nonmagnetized plasma examined in this study, manifests primarily as temporally broadened signals broadening, exhibiting strong frequency dependence with enhanced broadening at low frequencies. In the literature, this phenomenon is attributed to wave diffraction at a plasma scattering screen, which is implemented in the EMPI code through diffractive regime modeling. Only rays within a specific effective radius of the scattering screen can reach the observer. Because these rays follow varying geometric trajectories from different points within the effective circle, their arrival times differ, resulting in an extended signal duration that appears as a scattering tail in the waterfall plot. At lower frequencies, the effective radius increases, leading to a wider range of arrival times, and consequently, more pronounced signal broadening. This characteristic tail-like feature remains evident even after dedispersion, as illustrated in Figure 11. (The figure also reveals a rising wing preceding the main pulse flattening, which comes from the intrinsic pulse being a top-flat square wave, deliberately chosen to accentuate these effects, with both the scattering tail and the rising wing being influenced by the dedispersion process).
@@
4.1.5 Convergence Test
For completeness, systematic resolution tests across key output parameters are presented in this section to validate numerical convergence. As with any grid-based calculation, discretizing the domain introduces truncation errors that diminish with resolution refinement. While higher resolution generally improves accuracy by approaching the continuum solution, the convergence rate and threshold depend critically on system-specific physics.
The code under investigation produces two types of outputs – the simulated received signal and the plasma lensing gain – both of which are influenced by the resolution of the screen. The following demonstrates how these outputs converge to stable values as the resolution is progressively refined.